# When Are Bias-Free ReLU Networks Effectively Linear Networks?

### v2

@@ Appendix E Deep ReLU Network Learning Dynamics
$f({\bm{x}};{\bm{W}})=u{\bm{W}}_{L}\cdots{\bm{W}}_{2}\begin{bmatrix}{\bm{r}}_{1}^{+}\\ {\bm{0}}\end{bmatrix}{\bm{r}}^{\top}{\bm{x}}=\cdots=\frac{u^{L-1}}{(\sqrt{2})^{L-2}}{\bm{W}}_{L}\begin{bmatrix}{\bm{r}}_{L-1}^{+}\\ {\bm{0}}\end{bmatrix}{\bm{r}}^{\top}{\bm{x}}=\left(\frac{u}{\sqrt{2}}\right)^{L}{\bm{r}}^{\top}{\bm{x}}.$
@@
$f({\bm{x}};{\bm{W}})=u{\bm{W}}_{L}\cdots{\bm{W}}_{2}\begin{bmatrix}{\bm{0}}\\ {\bm{r}}_{1}^{-}\end{bmatrix}{\bm{r}}^{\top}{\bm{x}}=\cdots=\frac{u^{L-1}}{(\sqrt{2})^{L-2}}{\bm{W}}_{L}\begin{bmatrix}{\bm{0}}\\ {\bm{r}}_{L-1}^{-}\end{bmatrix}{\bm{r}}^{\top}{\bm{x}}=\left(\frac{u}{\sqrt{2}}\right)^{L}{\bm{r}}^{\top}{\bm{x}}.$
@@
Equation 16 is thus proved.
We now prove that under Equation 15 on the weights at time $t=0$, we have that $\forall\,t\geq 0$, Equation 15 remains valid. We assume that $\sigma^{\prime}(0)=0$. We substitute the low-rank weights defined in Equation 15 into the learning dynamics of deep bias-free ReLU networks and make simplifications. For the first layer,
@@
Appendices E, E and E can be written more compactly as Equation 17 if we substitute the weights back in. The dynamics of the deep ReLU network as in Equation 17 is the same as a deep linear network as in Equation 7 except for constant coefficients.
@@
After the weights have formed the low-rank structure specified in Equation 15, the norm of each layer $u$ and the rank-one direction of the first layer ${\bm{r}}$ evolve while ${\bm{r}}_{1},{\bm{r}}_{2},\cdots,{\bm{r}}_{L-1}$ stay fixed. Hence, under Equation 15 on the weights at time $t=0$, we have that $\forall\,t\geq 0$, Equation 15 remains valid.
@@ Appendix F Depth Separation
We visualize $g({\bm{x}})$ defined in Equation 9 below. This function cannot be expressed by any two-layer bias-free ReLU network because it is odd and nonlinear. Corollary 2 indicates that odd and nonlinear functions cannot be expressed by two-layer bias-free ReLU networks. It can be expressed by a three-layer bias-free ReLU network as written in Equation 9.
@@
Another relevant fact is that there is also depth separation between two-layer and deep ReLU networks with bias. One example is the pyramid function, $\sigma(1-\|{\bm{x}}\|_{1})$, which was studied in (Ongie et al., 2020, Example 4) and (Nacson et al., 2023, Proposition 2)
@@ Appendix H Implementation Details
Figure 1. The networks have width 100. The initialization scale $w_{\mathrm{init}}=10^{-2}$. The learning rate is $0.2$. The two-layer networks are trained 10000 epochs. The three-layer networks are trained 80000 epochs. The dataset is plotted in the figure. Empty black circles denote data points with $+1$ labels. Short black lines denote data points with $-1$ labels. The size of the datasets is 120.
@@
Figures 3 and 7. We use the same hyperparameters as Boursier et al. (2022). The network width is 60. The initialization scale $w_{\mathrm{init}}=10^{-6}$. The learning rate is $0.001$ for square loss and $0.004$ for logistic loss. The orthogonal input dataset contains two data points, i.e., $[-0.5,1],[2,1]$. The XOR input dataset contains four data points, i.e., $[0,1],[2,0],[0,-3],[-4,0]$.
@@
Figure 5. The network width is 100. The initialization scale $w_{\mathrm{init}}=10^{-3}$. The learning rate is $0.025$. The dataset contains six data points: $[1,1],[-1,-1],[1,-1],[-1,1],[-1,0],[1,\Delta y]$.

### v3

@@ Appendix E Deep ReLU Network Learning Dynamics
Refer to captionRefer to caption
Refer to captionRefer to caption
Figure 10: Same as Figure 5 but with deeper networks.
Proof of Proposition 12.
@@
$f({\bm{x}};{\bm{W}})=u{\bm{W}}_{L}\cdots{\bm{W}}_{2}\begin{bmatrix}{\bm{r}}_{1}^{+}\\ {\bm{0}}\end{bmatrix}{\bm{r}}^{\top}{\bm{x}}=\frac{u^{L-1}}{(\sqrt{2})^{L-2}}{\bm{W}}_{L}\begin{bmatrix}{\bm{r}}_{L-1}^{+}\\ {\bm{0}}\end{bmatrix}{\bm{r}}^{\top}{\bm{x}}=\left(\frac{u}{\sqrt{2}}\right)^{L}{\bm{r}}^{\top}{\bm{x}}.$
@@
$f({\bm{x}};{\bm{W}})=u{\bm{W}}_{L}\cdots{\bm{W}}_{2}\begin{bmatrix}{\bm{0}}\\ {\bm{r}}_{1}^{-}\end{bmatrix}{\bm{r}}^{\top}{\bm{x}}=\frac{u^{L-1}}{(\sqrt{2})^{L-2}}{\bm{W}}_{L}\begin{bmatrix}{\bm{0}}\\ {\bm{r}}_{L-1}^{-}\end{bmatrix}{\bm{r}}^{\top}{\bm{x}}=\left(\frac{u}{\sqrt{2}}\right)^{L}{\bm{r}}^{\top}{\bm{x}}.$
@@
Equation 19 is thus proven.
We now prove that under Equation 18 on the weights at time $t=t_{0}$, we have that $\forall\,t\geq t_{0}$, Equation 18 remains valid. We assume that $\sigma^{\prime}(0)=0$. We substitute the low-rank weights defined in Equation 18 into the learning dynamics of deep bias-free ReLU networks and make simplifications. For the first layer,
@@
Appendices E, E and E can be rewritten as Equation 20 if we substitute the weights back in. The dynamics of the deep ReLU network as in Equation 20 is the same as a deep linear network as in Equation 7 except for constant coefficients.
@@
After the weights have formed the low-rank structure specified in Equation 18, the norm of each layer $u$ and the rank-one direction of the first layer ${\bm{r}}$ evolve while ${\bm{r}}_{1},{\bm{r}}_{2},\cdots,{\bm{r}}_{L-1}$ stay fixed. Hence, under Equation 18 on the weights at time $t=t_{0}$, we have that $\forall\,t\geq t_{0}$, Equation 18 remains valid. ∎
We complement Figure 5 in the main text with Figure 10, which shows the converged weights in deeper bias-free linear and ReLU networks. Figures 5 and 10 are the empirical results that motivate us to make Conjecture 11.
@@ Appendix F Depth Separation
Another relevant fact is that there is also depth separation between two-layer and deep ReLU networks with bias. One example is the pyramid function, $\sigma(1-\|{\bm{x}}\|_{1})$, which was studied in (Ongie et al., 2020, Example 4) and (Nacson et al., 2023, Proposition 2).
@@ Appendix H Implementation Details
Figure 1. The networks have width 100. The initialization scale $w_{\mathrm{init}}=10^{-2}$. The learning rate is $0.2$. The two-layer networks are trained 10000 epochs. The three-layer networks are trained 80000 epochs. The dataset is plotted in the figure. The size of the datasets is 120.
@@
Figures 4 and 8. We use the same hyperparameters as Boursier et al. (2022). The network width is 60. The initialization scale $w_{\mathrm{init}}=10^{-6}$. The learning rate is $0.001$ for square loss and $0.004$ for logistic loss. The orthogonal input dataset contains two data points, i.e., $[-0.5,1],[2,1]$. The XOR input dataset contains four data points, i.e., $[0,1],[2,0],[0,-3],[-4,0]$.
@@
Figure 6. The network width is 100. The initialization scale $w_{\mathrm{init}}=10^{-3}$. The learning rate is $0.025$. The dataset contains six data points: $[1,1],[-1,-1],[1,-1],[-1,1],[-1,0],[1,\delta]$.